### v4

@@ 5.1.2 Implementation Details
We implement the method using Tensorflow Abadi et al. (2016) and Pytorch Paszke et al. (2019). The feature embedding size is set as 16 for all methods. For a fair comparison, we set three layers for AutoInt, FiGNN, and also GraphFM. There are two attention heads in AutoInt and GraphFM. Implementation of the other compared baselines follows Song et al. (2019) and  Tian et al. (2023). The optimal hyper-parameters are found by grid search. $\left[m_{1},m_{2},m_{3}\right]$ are set as $\left[39,20,5\right]$, $\left[23,10,2\right]$, and $\left[7,4,2\right]$ for Criteo, Avazu and MovieLens-1M dataset respectively. We use Adam Kingma and Ba (2014) to optimize all these models. The experiments were conducted over a server equipped with 8 NVIDIA Titan X GPUs.
@@ 5.2.1 Evaluation of Effectiveness
The performance comparison of these methods on three datasets is presented in Table 2, from which we have the following observations. Our proposed GraphFM achieves best performance among all these four classes of methods on three datasets. The performance improvement of GraphFM compared with the three classes of methods (A, B, C) is especially significant, above $\mathbf{0.01}$-level. The aggregation-based methods including InterHAt, AutoInt, Fi-GNN and our GraphFM consistently outperform the other three classes of models, which demonstrates the strength of the aggregation strategy in capturing high-order relations. Compared with the strong aggregation-based baselines AutoInt and Fi-GNN, GraphFM still advances the performance by a large margin, especially on MovieLens-1M dataset. The performance improvement on the other two datasets are also at $\mathbf{0.001}$-level, which can be regarded as significant for CTR prediction task Cheng et al. (2016); Guo et al. (2017); Song et al. (2019). Such improvement can be attributed to its combination with FM, which introduces feature interactions operations, and also the interaction selection mechanism, which selects and models only the beneficial feature interactions. GraphFM outperforms the compared baselines by the largest margin on MovieLens-1M dataset, whose feature size is smallest among the three datasets. I suppose this is because the feature embedding size is not large enough for the other two datasets.
@@
Figure 2: The performance comparison of GraphFM with different components on Criteo and MovieLens-1M datasets. Further analysis is provided in Section 5.3.
@@ 5.3 Ablation Studies
  • GraphFM(-S): interaction selection is the first component in each layer of GraphFM, which selects only the beneficial feature interactions and treat them as edges. As a consequence, we can model only these beneficial interactions with the next interaction aggregation component. To check the necessity of this component, we remove this components, so that all pair of feature interactions are modeled as a fully-connected graph.
  • GraphFM(-I): in the interaction aggregation componnet, we aggregate the feature interactions instead of the neighbors’ features as in the standard GNNs. To check its rationality, we test the performance of directly aggregating neighborhood features instead of the feature interactions with them.
@@
  • GraphFM(-$K_{i}$): before obtaining the final representation of the feature, we concatenate and average the feature representation $e_{k}$ output from layer $k$. In order to study the degree of contribution of the features learned at different layers to the results, we use the feature representation of $k$-th layer for direct prediction, and there are $K=3$ layers.
The performances of GraphFM and these Four variants are shown in Fig. 2. We observe that GraphFM outperforms all the ablative methods, which proves the necessity of all these components in our model. The performance of GraphFM(-M) suffers from a sharp drop compared with GraphFM, proving that it is necessary to transform and aggregate the feature interactions in multiple semantic subspaces to accommodate the polysemy. Note that although we did not present the statistics here, we also tested the influence of number of attention heads $H$. The performance of using only one head, i.e., GraphFM(-M), is worse than that of using two, and more attention heads don not lead to improvement of performance but introduce much higher time and space complexity. GraphFM(-I) does not perform well either. This is reasonable, as without the interaction between features, the neighborhood aggregation operation will only make the neighboring features similar. As a consequence, no feature interactions are guaranteed to be captured. This interaction part is also the most significant difference between GraphFM and GNN, and the resulting gap in term of performance indicates that GraphFM is able to leverage the strength of FM to solve the drawbacks of GNN in modeling feature interactions. GraphFM(-S) achieves slightly worse performance than GraphFM, demonstrating that selecting and modeling only the beneficial interactions instead of all of them can avoid the noise and make it easier to train the model. The GraphFM(-$K_{i}$) results show that each layer can learn features that benefit the results. Although different features can have a positive effect on the task prediction, the difference in effect is not large and the prediction results are worse than those obtained by combining all features, demonstrating the need to merge the features of the $k$-th layers.
@@ 5.4 Study of Neighborhood Sampled Size
The number of selected neighbors for each feature node is an important hyper-parameter, which controls the number of features that each feature interacts with. We thus investigate how the neighborhood sampled size effects the model performance. As the total search space is too large, we here only show the performance of our model with $K=3,m_{1}=n$ and varying value of $m_{2}$ and $m_{3}$. The results on three datasets are summarized in Fig. 3.
On Criteo dataset, there are totally 39 feature fields. We found that our model achieves best performance with $m_{1}=39,m_{2}\times m_{3}=100$. The performances vary in the range of [0.8084, 0.8091], which proves that our model is quite robust, and not very sensitive to the size of neighborhood sampled size. On Avazu dataset, the model performance peaks with $m_{1}=23,m_{2}=10,m_{3}=2$ or $m_{2}=15,m_{3}=4$. On MovieLens-1M dataset, the model performance peaks when $m_{2}\times m_{3}$ at around 9. We also found a diminishing trend for sampling larger or smaller neighborhoods. In a word, the optimal neighborhood sampled sizes depend on the sizes of datasets.
@@
Figure 4: Heat maps of estimated edge weights of correctly predicted instance (a) and wrongly predicted instance (b) on MovieLens-1M dataset, where positive edge weights indicate beneficial feature interactions. The axises represent feature fields (Gender, Age, Occupation, Zipcode, ReleaseTime, WatchTime, Genre).
5.5 Visualization of Interaction Graph
@@
Since the features along with selected beneficial feature interactions are treated as a graph, it can provide human readable interpretations on the prediction. Here we visualize heat maps of estimated edge weights of two cherry-pick instances on MovieLens-1M dataset in Fig. 4. We show the measured edge weights of each instance in the three layers, which selects the order-2, order-3, and order-4 feature interactions respectively. The positive edge weights indicate how beneficial feature interactions are. We set $S_{1}=7,S_{2}=4,S_{3}=2$, which means that we only keep 7, 4, and 2 pairs of beneficial order-2, order-3, and order-4 feature interactions respectively. Therefore, there are only 7, 4, and 2 interacted feature fields for each feature fields each row for heat maps of order-2, order-3, and order-4. The axises represent feature fields (Gender, Age, Occupation, Zipcode, ReleaseTime, WatchTime, Genre). Gender, Age, Occupation and Zipcode are users’ demographic information. Note Zipcode indicates the users’ place of residence. textitReleaseTime and Gender are the movie information. WatchTime (Timestamp) represents the time when users watched the movies.
@@
From the two instances in Fig. 4, we can obtain some following interesting observations. We find that in the first layer, which models the second order feature interactions, these feature fields are hard to distinguish when selecting the beneficial interactions. This suggests that almost all the second-order feature interactions are useful, which also why we sample all of them in the first layer, i.e., $m_{1}=n$. Except that the diagonal elements are with the smallest values, which suggests that our designed interaction selection mechanism can classify the redundant self-interacted feature interactions, even though we keep and model all pairs of feature interactions. The selected feature interactions of order-3 and order-4 are mostly not overlapped in the correctly predicted instance (a). In instance (a), our model selects relevant feature fields (Gender, Age, ReleaseTime, WatchTime) for Genre in order-3, while selects the other two feature fields (Occupation, Gender) in order-4. However, in the wrongly predicted instances (b), the feature interactions of order-3 and order-4 are mostly not overlapped.
@@
This proves that our model can indeed select meaningful feature combination and model feature interactions of increasing orders with multiple layers in most cases, rather than select the redundant feature combinations of same feature fields. We can also find some meaningful feature combinations in common cases. For examples, Gender are usually relevant with feature fields Age, occupation, and WatchTime, while Age are usually relevant with feature fields Gender, WatchTime, and Genre. This provides some rationales for the model prediction.
@@ 6 Conclusion and Future Work
In this work, we disclose the relationship between FM and GNN, and seamlessly combine them to propose a novel model GraphFM for feature interaction learning. The proposed model leverages the strength of FM and GNN and also solve their respective drawbacks. At each layer of GraphFM, we select the beneficial feature interactions and treat them as edges in a graph. Then we utilize a neighborhood/interaction aggregation operation to encode the interactions into feature representations. By design, the highest order of feature interaction increases at each layer and is determined by layer depth, and thus the feature interactions of order up to the highest can be learned. GraphFM models high-order feature interactions in an explicit manner, and can generate human readable explanations on model outcome. Experimental results show that GraphFM outperforms the state-of-the-art baselines by a large margin. In addition, we conduct extensive experiments to analyze how the highest order of feature interactions and number of modeled feature interactions influence the model performance, which can help us gain deeper insight into feature interaction modeling. In the future, we aim to investigate whether the proposed method can also benefit graph representation learning, and graph/node classification tasks.
@@ Acknowledgements
We would like to thank the anonymous reviewers for their valuable comments and suggestions allowing to improve the quality of this paper. This work is sponsored by National Science Foundation of China (62141608).

### v5

@@ 5.1.2 Implementation Details
We implement the method using TensorFlow Abadi et al. (2016) and Pytorch Paszke et al. (2019). The feature embedding size is set as 16 for all methods. For a fair comparison, we set three layers for AutoInt, FiGNN, and GraphFM. There are two attention heads in AutoInt and GraphFM. The implementation of the other compared baselines follows Song et al. (2019) and  Tian et al. (2023). The optimal hyper-parameters are found via a grid search. $\left[m_{1},m_{2},m_{3}\right]$ are set as $\left[39,20,5\right]$, $\left[23,10,2\right]$, and $\left[7,4,2\right]$ for the Criteo, Avazu and MovieLens-1M datasets respectively. We use Adam Kingma and Ba (2014) to optimize all these models. The experiments were conducted over a server equipped with 8 NVIDIA Titan X GPUs.
@@ 5.2.1 Evaluation of Effectiveness
The performance comparison of these methods on three datasets is presented in Table 2, from which we have the following observations. Our proposed GraphFM achieves the best performance among all four classes of methods on three datasets. The performance improvement of GraphFM compared with the three classes of methods (A, B, C) is especially significant, above the $\mathbf{0.01}$-level. The aggregation-based methods including InterHAt, AutoInt, Fi-GNN and our GraphFM consistently outperform the other three classes of models, which demonstrates the strength of the aggregation strategy in capturing high-order relations. Compared with the strong aggregation-based baselines AutoInt and Fi-GNN, GraphFM still achieves large improvements in performance, especially on the MovieLens-1M dataset. The performance improvement on the other two datasets is also at the $\mathbf{0.001}$-level, which can be regarded as significant for the CTR prediction task Cheng et al. (2016); Guo et al. (2017); Song et al. (2019); Zhang et al. (2022); Wu et al. (2020); Xu et al. (2021). This improvement can be attributed to its combination with FM, which introduces feature interaction operations, and the interaction selection mechanism, which selects and models only the beneficial feature interactions. GraphFM outperforms the compared baselines by the largest margin on the MovieLens-1M dataset, whose feature size is the smallest among the three datasets. This is likely because the feature embedding size is not large enough for the other two datasets.
@@
Figure 2: Performance comparison of GraphFM with different components on the Criteo and MovieLens-1M datasets. Further analysis is provided in Section 5.3.
@@ 5.3 Ablation Studies
  • GraphFM(-S): interaction selection is the first component in each layer of GraphFM, which selects only the beneficial feature interactions and treats them as edges. As a consequence, we can model only these beneficial interactions with the next interaction aggregation component. To check the necessity of this component, we remove these components, so that all pairs of feature interactions are modelled as a fully-connected graph.
  • GraphFM(-I): In the interaction aggregation component, we aggregate the feature interactions instead of the neighbors’ features, as in the standard GNNs. To check its rationality, we test the performance of directly aggregating neighborhood features instead of the feature interactions with them.
@@
  • GraphFM(-$K_{i}$): Before obtaining the final representation of the feature, we concatenate and average the feature representation $e_{k}$ output from layer $k$. To study the degree of contribution of the features learned at different layers to the results, we use the feature representation of the $k$-th layer for direct prediction, and there are $K=3$ layers.
The performances of GraphFM and these four variants are shown in Fig. 2. We observe that GraphFM outperforms all the ablative methods, which proves the necessity of all these components in our model. The performance of GraphFM(-M) suffers from a sharp decrease compared with that of GraphFM, proving that it is necessary to transform and aggregate the feature interactions in multiple semantic subspaces to accommodate polysemy. Note that although we did not present the statistics here, we also tested the influence of the number of attention heads $H$. The performance of using only one head, i.e., GraphFM(-M), is worse than that of using two, and more attention heads do not lead to improvement in performance but introduce much greater time and space complexity. GraphFM(-I) does not perform well either. This is reasonable, as without the interaction between features, neighborhood aggregation operation will only make neighboring features similar. As a consequence, no feature interactions are guaranteed to be captured. This interaction is also the most significant difference between GraphFM and GNN, and the resulting difference in terms of performance indicates that GraphFM is able to leverage the strength of FM to overcome the drawbacks of GNN in modelling feature interactions. GraphFM(-S) achieves slightly worse performance than GraphFM, demonstrating that selecting and modelling only the beneficial interactions instead of all of them can avoid noise and make it easier to train the model. The GraphFM(-$K_{i}$) results show that each layer can learn features that benefit the results. Although different features can have a positive effect on task prediction, the difference in effect is not large, and the prediction results are worse than those obtained by combining all features, demonstrating the need to merge the features of the $k$-th layers.
@@ 5.4 Study of Neighborhood Sampled Size
The number of selected neighbors for each feature node is an important hyper-parameter that controls the number of features with which each feature interacts. We thus investigate how the neighborhood sample size affects the model performance. As the total search space is too large, we only show the performance of our model with $K=3,m_{1}=n$ and varying values of $m_{2}$ and $m_{3}$. The results on three datasets are summarized in Fig. 3.
On the Criteo dataset, there are a total of 39 feature fields. We found that our model achieves the best performance with $m_{1}=39$, and $m_{2}\times m_{3}=100$. The performances vary in the range of [0.8084, 0.8091], which proves that our model is quite robust, and not very sensitive to the size of the neighborhood sampled. On the Avazu dataset, the model performance peaks with $m_{1}=23,m_{2}=10,m_{3}=2$ or $m_{2}=15,m_{3}=4$. On the MovieLens-1M dataset, the model performance peaks when $m_{2}\times m_{3}$ is approximately 9. We also found a diminishing trend for sampling larger or smaller neighborhoods. In other words, the optimal neighborhood sample size depends on the dataset size.
@@
Figure 4: Heat maps of the estimated edge weights of correctly predicted instances (a) and incorrectly predicted instances (b) on the MovieLens-1M dataset, where positive edge weights indicate beneficial feature interactions. The axes represent feature fields (Gender, Age, Occupation, Zipcode, ReleaseTime, WatchTime, Genre).
5.5 Visualization of the Interaction Graph
@@
Since the features along with selected beneficial feature interactions are treated as a graph, they can provide human readable interpretations of the prediction. Here, we visualize heat maps of the estimated edge weights of two cherry-pick instances on the MovieLens-1M dataset in Fig. 4. We show the measured edge weights of each instance in the three layers, which select the order-2, order-3, and order-4 feature interactions. The positive edge weights indicate how beneficial feature interactions are. We set $S_{1}=7$, $S_{2}=4$, and $S_{3}=2$, which means that we only retain 7, 4, and 2 pairs of beneficial order-2, order-3, and order-4 feature interactions respectively. Therefore, there are only 7, 4, and 2 interaction feature fields for each feature field in each row for heat maps of order-2, order-3, and order-4, respectively. The axes represent feature fields (Gender, Age, Occupation, Zipcode, ReleaseTime, WatchTime, Genre). Gender, Age, Occupation and Zipcode are users’ demographic information. Note that Zipcode indicates the users’ place of residence. textitReleaseTime and Gender are the movie information. WatchTime (Timestamp) represents the time when users watched the movies.
@@
From the two instances in Fig. 4, we can obtain the following interesting observations. We find that in the first layer, which models the second order feature interactions, these feature fields are difficult to distinguish when selecting the beneficial interactions. This suggests that almost all the second-order feature interactions are useful, which is why we sample all of them in the first layer, i.e., $m_{1}=n$, except that the diagonal elements have the smallest values, which suggests that our designed interaction selection mechanism can classify the redundant self-interacting feature interactions, even though we keep and model all pairs of feature interactions. The selected feature interactions of order-3 and order-4 mostly do not overlap in the correctly predicted instance (a). In instance (a), our model selects relevant feature fields (Gender, Age, ReleaseTime, and WatchTime) for Genre in order-3 but selects the other two feature fields (Occupation and Gender) in order-4. However, in the wrongly predicted instances (b), the feature interactions of order-3 and order-4 mostly do not overlap.
@@
This proves that our model can indeed select meaningful feature combinations and model feature interactions of increasing orders with multiple layers in most cases, rather than selecting the redundant feature combinations of the same feature fields. We can also find some meaningful feature combinations in common cases. For example, Gender is usually relevant to the feature fields Age, occupation, and WatchTime, while Age is usually relevant to the feature fields Gender, WatchTime, and Genre. This provides some rationale for the model prediction.
@@ 6 Conclusion and Future Work
In this work, we disclose the relationship between FMs and GNNs, and seamlessly combine them to propose a novel model GraphFM for feature interaction learning. The proposed model leverages the strengths of FMs and GNNs and solves their respective drawbacks. At each layer of GraphFM, we select the beneficial feature interactions and treat them as edges in a graph. Then, we utilize a neighborhood/interaction aggregation operation to encode the interactions into feature representations. By design, the highest order of feature interaction increases at each layer and is determined by layer depth; thus, the feature interactions of order up to the highest can be learned. GraphFM models high-order feature interactions in an explicit manner, and can generate human readable explanations of outcomes. The experimental results show that GraphFM outperforms the state-of-the-art baselines by a large margin. In addition, we conduct extensive experiments to analyse how the highest order of feature interactions and the number of modelled feature interactions influence model performance, which can help us gain deeper insight into feature interaction modelling. In the future, we aim to investigate whether the proposed method can also benefit graph representation learning, and graph/node classification tasks.
@@ Acknowledgements
We would like to thank the anonymous reviewers for their valuable comments and suggestions, allowing us to improve the quality of this paper. This work is sponsored by the National Science Foundation of China (62141608).